alones (alones):

Which one would be the right one ;3

alones (alones):

\[ \sqrt{20 * √5 } \]

alones (alones):

Will it be 10 or √10

OpenStudy (anonymous):

\[\sqrt20=\sqrt5 * \sqrt4\] So \[\sqrt 20(\sqrt5)\] is equal to: \[\sqrt5*\sqrt5*\sqrt4\] and \[\sqrt5*\sqrt5=5\]\[\sqrt4=2\] 5*2=10

alones (alones):

Oh okay thanks ;3

Directrix (directrix):

@AloneS I think that this answer is not correct.

alones (alones):

Oh

Directrix (directrix):

I worked it by hand and then decided to let the "Wolf" have a say.

Directrix (directrix):

Question: Would you check and see if you wrote the problem correctly?

Directrix (directrix):

I think this might be what the original problem is: